OpenStudy (anonymous):

2y=5x+11 Find 3 Ordered Pairs ,

OpenStudy (anonymous):

Just put some values into x and you will get y values... Ex: put x=1, you will get y=8 So, one ordered pair is (1,8)

OpenStudy (mayankdevnani):

value of x=0

OpenStudy (mayankdevnani):

2y=5*0+11

OpenStudy (mayankdevnani):

y=11/2

OpenStudy (mayankdevnani):

first order pair (0,11/2)

OpenStudy (anonymous):

0,5

OpenStudy (mayankdevnani):

value of x=1

OpenStudy (mayankdevnani):

2y=5+11 2y=17 y=17/2

OpenStudy (anonymous):

Thanks I Know How To Do It

OpenStudy (mayankdevnani):

second order pair (1,17/2)

OpenStudy (mayankdevnani):

ok welcome